Denver Springs predicts that this year's sales will total $7,500,000. The selling price for their software...

  1. Denver Springs predicts that this year's sales will total $7,500,000. The selling price for their software is $135 per unit. Variable costs amount to $80 per unit. The firm also has $6,000,000 in bonds outstanding with a coupon interest rate of 11%. Net income is projected to be $139,500. The firm's marginal tax rate is 34%. What is the firm's total fixed costs?
Answer #1

Answer :

Earning Before Interest and Tax = Sales - Variable Cost - Fixed Cost.........(1)

Calulation of Earning Before Interest and Tax

Earning Before Tax = Earning Before Interest and Tax - Interest

As net income is after tax so it represesnts 66% (100% - 34%) of Earning Before Tax

Therefore Earning Before Tax = 139500 / 0.66 = 211363.636363

211363.636363 = Earning Before Interest and Tax - (6,000,000 * 11%)

Earning Before Interest and Tax = 211,363.636363 + 660,000

Earning Before Interest and Tax = 871,363.636363

Putting Earning Before Interest and Tax in Equation (1)

Earning Before Interest and Tax = 7,500,000 - [(7,500,000 / 135) * 80] - Fixed Cost

871363.636363 = 7,500,000 - 4,444,444.44444 - Fixed Cost

==>Fixed Cost = 7,500,000 - 4,444,444.44 - 871,363.636363

= 2,184,191.9192 or 2,184,191.92